# A Compute-Matched Re-Evaluation of TroVE on MATH

### v1

@@ 1 Introduction
However, Berlot-Attwell et al. (2024) raise the question whether the TroVE mechanism actually increases performance as they show that tools are either trivial (already contained in the knowledge of the LLM) or never reused: An ablation study in a subset of MATH shows that the removal of the Import mode does not negatively affect performance. Although the authors hypothesize TroVE’s benefit may come from self-consistency or the self-correctness mechanism, the question of what makes TroVE better remains an open question.
@@ 4.1.1 Matching Computational Budget
As depicted in Table 1, the compute-matched Primitive baseline now, instead of reported , achieves in MATH, outperforming the original 22% reported by TroVE. Among the different categories, the performance differs at most in accuracy between Primitive and TroVE with Primitive performing better in algebra, geometry, intermediate algebra, number and prealgebra. When analyzing the performance across different computational budgets, the results suggest that the reported results of Primitive in the original work were achieved by setting to instead of (see Appendix B.1, Table 5). Thus, contrary to the claim of the authors, TroVE appears to benefit primarily from repeated sampling rather than toolbox learning.
@@ 6 Conclusion
The findings of this study indicate that the primary advantage of the TroVE approach comes less from the use of an incrementally assembled toolbox of abstractions and more from the increased probability of finding the correct solutions through repeated sampling. When controlling for the total number of generated solutions, the baseline approach that simply samples more candidate programs using its inherent domain knowledge performs comparable to TroVE. In the MATH domain, we confirm that library learning does not bring a significant advantage over this simple baseline. However, our experiments show that TroVE’s prompting mechanism does lead to an increased hypothesis space. Under a stronger selection mechanism, this may lead to a small benefit over Primitive, which exists in our experiments but under majority selection we demonstrate it is not significant. More importantly, a driving factor is the choice of the selection mechanism. Although TroVE and Primitive often create a correct candidate, the selection mechanism often does not select it, as shown by the difference between oracle selection and majority voting for . Improving the current selection mechanism would lead to a much higher benefit than the one currently brought from TroVE over the baseline.
In conclusion, while our results cast doubt on whether tool-based methods represent an alternative to intensive sampling in MATH, we remain optimistic about the long-term potential of systematic abstraction learning for LLMs in other domains. Recent work on agentic tasks (Wang et al., 2025; Zheng et al., 2025) has shown that tools can effectively compress past experiences that can then be reused to improve efficiency and solving harder long horizon challenges. By deepening understanding on when and how library learning can assist LLMs, future research can more effectively utilize toolboxes to achieve improvements in complex problem solving.
7 Impact Statement
@@ Appendix A Experimental Setup
The experiments were run on a single NVIDIA RTX A6000, 2 CPUs, and 48GB RAM.

### v2

@@ 1 Introduction
However, Berlot-Attwell et al. (2024) raise the question whether the TroVE mechanism actually increases performance as they show that tools are either trivial (already contained in the knowledge of the LLM) or never reused (see Appendix B): An ablation study in a subset of MATH shows that the removal of the Import mode does not negatively affect performance. Although the authors hypothesize TroVE’s benefit may come from self-consistency or the self-correctness mechanism, the question of what makes TroVE better remains an open question.
@@ 4.1.1 Matching Computational Budget
As depicted in Table 1, the compute-matched Primitive baseline now, instead of reported , achieves in MATH, outperforming the original 22% reported by TroVE. Among the different categories, the performance differs at most in accuracy between Primitive and TroVE with Primitive performing better in algebra, geometry, intermediate algebra, number and prealgebra. When analyzing the performance across different computational budgets, the results suggest that the reported results of Primitive in the original work were obtained by setting to a much smaller value than (see Appendix B.1, Table 6). Thus, contrary to the claim of the authors, TroVE appears to benefit primarily from repeated sampling rather than toolbox learning. Preliminary results (see Appendix B.7) suggest that this observation also applies to domains beyond MATH, such as TabMWP (Lu et al., 2022), WTQ (Pasupat & Liang, 2015), and HiTab (Cheng et al., 2021).
@@ 6 Conclusion
The findings of this study indicate that the primary advantage of the TroVE approach comes less from the use of an incrementally assembled toolbox of abstractions and more from the increased probability of finding the correct solutions through repeated sampling. When controlling for the total number of generated solutions, the baseline approach that simply samples more candidate programs using its inherent domain knowledge performs comparable to TroVE. In the MATH domain, we confirm that library learning does not bring a significant advantage over this simple baseline. Preliminary results indicate that this observation also applies to further domains (TabMWP, WTQ and HiTab). However, our experiments show that TroVE’s prompting mechanism does lead to an increased hypothesis space. Under a stronger selection mechanism, this may lead to a small benefit over Primitive, which exists in our experiments but under majority selection we demonstrate it is not significant. More importantly, a driving factor is the choice of the selection mechanism. Although TroVE and Primitive often create a correct candidate, the selection mechanism often does not select it, as shown by the difference between oracle selection and majority voting for . Improving the current selection mechanism would lead to a much higher benefit than the one currently brought from TroVE over the baseline.
In conclusion, while our results cast doubt on whether tool-based methods represent an alternative to intensive sampling in MATH, we remain optimistic about the long-term potential of systematic abstraction learning for LLMs in other problem areas. Recent work on agentic tasks (Wang et al., 2025; Zheng et al., 2025) has shown that tools can effectively compress past experiences that can then be reused to improve efficiency and solving harder long horizon challenges. By deepening understanding on when and how library learning can assist LLMs, future research can more effectively utilize toolboxes to achieve improvements in complex problem solving.
7 Limitations
A main limitation of our work is that our re-evaluation is primarily focused on the MATH dataset. Although preliminary studies suggest that the observation also applies to TabMWP, WTQ and HiTab, they need to be further analyzed.
Second, we evaluate both approaches exclusively using the CodeLLaMa2-7b-Instruct model. Although this aligns with the original evaluation of TroVE, further experiments involving a broader range of LLMs would help validate the generality of our conclusions.
Third, we assess TroVE’s upper-bound performance under an oracle selection mechanism, which assumes access to the correct answer within the candidate set. While this provides insights into the potential benefits of prompt diversity, it is not applicable in practice. We do not propose or evaluate any realistic selection strategy capable of consistently leveraging this potential, leaving this an interesting direction for future work.
Moreover, our analysis is limited to quantitative performance. We do not perform a qualitative investigation into the nature of the tools generated or their reuse, which lies outside the scope of this work. For a detailed examination of tool reuse in TroVE, we refer the reader to Berlot-Attwell et al. (2024) and Appendix B.6.
Finally, we use the number of LLM calls as a proxy for computational cost when comparing the two approaches. However, this approximation does not imply equal runtime performance. In practice, even with the same number of LLM calls, the Primitive baseline exhibits faster execution, likely due to more efficient batching and reduced overhead from its simpler structure.
Impact Statement
@@ 7 Limitations
Acknowledgements
This work is partially supported by the Federal Ministry of Education and Research (BMBF) and by the Federal Ministry for Economic Affairs and Energy of Germany (BMWE).
@@ Appendix A Experimental Setup
The experiments were run on a single NVIDIA RTX A6000, 2 CPUs, and 48GB RAM. Our code and data are available at https://github.com/Tsesterh/TroVE_Compute_Matched.
As shown in Table 5, although we match Primitive in terms of LLM call budget, the baseline method runs consistently faster than TroVE. We propose that this is due to Primitive batching 15 LLM calls, whereas TroVE iteratively calls the LLM 5 times per mode and performs additional postprocessing, i.e., periodical toolbox trimming.
@@
| Method | Algebra | Geometry | Counting | Intermediate | Number | Prealgebra | Precalculus | MATH |
| Primitive | 25 | 5 | 7 | 14 | 12 | 17 | 4 | 84 |
| TroVE | 38 | 12 | 12 | 25 | 26 | 28 | 7 | 148 |
@@ B.5 Distribution among MATH difficulties
B.6 Tool Reuse
As this work complements the work by Berlot-Attwell et al. (2024), it focuses solely on the quantitative re-evaluation of TroVE. However, Berlot-Attwell et al. (2024) perform a deeper analysis on tool reuse on MATH, across all seven categories. Their findings show that TroVE only learns functions for 3 of the 7 areas — counting, number, and prealgebra — resulting in a total of just 15 learned functions. No functions are learned for algebra, geometry, intermediate algebra, or pre-calculus. Moreover, only two of these functions are ever reused in a correct solution: is_perfect_square(n) once and is_prime(num) twice. With just three reuses across 3,201 test questions, the authors conclude that TroVE’s performance gains are unlikely to stem from tool reuse, which is elaborated in our work by showing that the benefit simply comes from increased compute.
B.7 Analysis on Other Domains
To extend our evaluation to additional domains, we test the Primitive baseline on three more datasets: TabMWP (Lu et al., 2022), WTQ (Pasupat & Liang, 2015), and HiTab (Cheng et al., 2021). We run experiments using a single random seed and three different computational budgets (). As shown in Figure 11, when compute-matched with , Primitive performs much more similarly to TroVE than suggested by the originally reported results. Specifically, accuracy improves by 3% on TabMWP, 5% on WTQ, and 6% on HiTab.
To validate these trends more robustly, we propose evaluating across multiple random seeds in future work. Nonetheless, these preliminary results suggest that our core observations from MATH may generalize to other domains as well, indicating that the performance gap between TroVE and Primitive may be marginal when computational budgets are matched.
| Method | TabMWP | WTQ | HiTab |
| Primitive | 0.36 | 0.18 | 0.08 |
| Primitive | 0.42 | 0.23 | 0.14 |
| Primitive | 0.46 | 0.25 | 0.15 |
| Primitive (Wang et al., 2024b) | 0.43 | 0.20 | 0.09 |
| TroVE (Wang et al., 2024b) | 0.47 | 0.21 | 0.18 |